OpenStudy (anonymous):

Please help me out. :) How many excess electrons must be placed on each of two spheres 3cm apart if the force of repulsion between the spheres is 10^-19 N?

ganeshie8 (ganeshie8):

use coulomb's law ?

ganeshie8 (ganeshie8):

say you need to put `q` amount of charge on each of the spheres : \[F = 9\times 10^9 ~ \dfrac{q\times q}{r^2}\] plugin the given values and solve \(q\)

OpenStudy (anonymous):

ahh.. okay.. :) wait..

ganeshie8 (ganeshie8):

take ur time, make sure you convert the distance to appropriate unit : meter

OpenStudy (anonymous):

sure.. thanks for reminding me. :)

OpenStudy (anonymous):

the value of q is 1 x 10^-16.. I'll convert it to e- .. so my final answer is 624.22 e- :) am I right? :)

ganeshie8 (ganeshie8):

wolfram says \(\large q = 3 \times 10^{-16}\) ?